Classify each system. Give the number of solutions.
{y=-x+5
{x+y=5

Answer by rapaljer(4671)

{y=-x+5
{x+y=5

In the first equation, add +X to each side, and you get x+y =5. This is exactly the same as the second equation, so the solution is the entire line. The system is said to be "dependent," and there are infinitely many solutions.

R^2